OpenStudy (anonymous):

Hi, I really need help to solve this question - I don't understand it. I need to solve it by hand and not a calculator or spreadsheet. Thanks! Determine an interest rate less than 15%, a period of investment greater than two years, and a regular payment that will result in the total amount of interest you earn being equal to the total amount of money you put in? (for example, under what conditions will you have a future value of $1 000 000, having earned $500 000 interest?)

OpenStudy (is3535):

Want [(1+r)^n - 1] / r = 2 where r < 15% and n > 2 years. Rule of 72 works. Try 8% and 9 years.

OpenStudy (anonymous):

I don't understand may you explain? Thanks I appreciate you helping me

OpenStudy (is3535):

(1+r)^t = 2 or t*ln(1+r) = ln 2. or t = ln(2)/ln(1+r) Now just start plugging in values of r and you will get some value of t. Then graph them.

OpenStudy (anonymous):

I saw those answers on yahoo answers but got confused hence i am here lol but thanks

OpenStudy (anonymous):

woops I got it sorry I was half alseep!